# Quantum multiple gray scale images encryption scheme in the bit plane representation model

### 01 Jan 2024 v1

@@
Quantum multiple gray scale images encryption scheme in the bit plane representation model
Claire I. Levaillant
clairelevaillant@yahoo.fr
@@ 1 Introduction
Chaotic systems are commonly used as diffusion tools in quantum image encryption methods. The idea to use chaos in encryption algorithms originates in [13]. Chaotic functions have the characteristics of generating completely different sequences through relatively small changes in the tuning parameters or in the initial conditions. One of the best-studied such functions is the logistic function which has been used to model biological populations. However, the parameter range for chaotic behavior is relatively small and for some isolated parameter values, the behavior is not chaotic. A general way to enlarge the parameter range of chaos and in fact make this range infinite is to use a sine chaotification model composed with another given chaotic map, see [7]. This idea was used by the authors in [11], applied to the $1D$ logistic map. In the present paper we also apply this idea with the base chaotic map being the $2D$ Hénon map. For each scrambled image, we use a different control parameter, while the two initial conditions depend on the multi-image plaintext. For a given scrambled image, the two pseudo-random sequences obtained, of length the side size of each image, are sorted by increasing order (we impose the elements of each sequence to be distinct with one another). The resulting two pseudo-permutations are then used in the following manner. Each element of the first (resp second) pseudo-random sequence gets applied a Chebyshev polynomial with index issued from the second (resp first) pseudo-random permutation and the resulting two values get multiplied together yielding an element of the interval $[-1,1]$. For each pixel position we thus obtain an element of this interval. We then multiply it by a power of ten depending on the image and take the integer part modulo the $\lceil log_{2}L\rceil$-th power of two, where $L$ is the number of bit planes. Each integer is then converted into a sequence of bits of length $L$, namely the secret keys, that are ready to be XOR-ed with the bit values of each bit plane.
@@ 2 The BRQMI: a bit-plane representation for quantum multiple images
where $P_{lxy}$ is the bit value in the $l$-th bit plane at pixel $(x,y)$. We now introduce the new model BRQMI used for our encryption/decryption scheme. Our BRQMI is a combination of BRQI and of the quantum representation model for multiple images QRMMI of [22]. With respect to BRQI, $\lceil log_{2}M\rceil$ qubits are added in order to encode the $m$-th image, with $10\leq m\leq M-$ 1 and $M$ the total number of images to be encrypted. Denoting by $L$ the number of bit planes, the quantum image is the superposition:
@@
$P_{mlxy}$ is the bit value in the $l$-th bit plane of the $m$-th image at pixel $(x,y)$. The 00-th bit plane denotes the least significant bit plane and the $(L-1)$-th bit plane is the most significant bit plane. With this representation, it is understood that we set:
@@
Moreover, for the sake of our encryption/decryption scheme, because as part of the two stage scrambling process we will apply a quantum baker map in the $(m,l)$ plane, we impose that $M=L$. This will be assumed throughout the rest of the paper. As far as the conversion of the classical multiple images into the quantum multiple images, it does translate in the following way. On one hand, if $M>L$, that is there are more images than bit planes, then we add more bit planes filled with bit values 00. On the other hand, if $M<L$, that is there are less images than bit planes, then we add $L-M$ blank images. It has to be determined in both cases how this would affect the security of the protocol. We note in particular that in the case when there are more bit planes than images, the attacker has no way of deciding on the number of images that are encrypted. This is rather a factor increasing the security of the scheme.
@@ 3 The quantum baker map used for scrambling
[…]
@@
Proof. Suppose $N_{i-1}\leq\frac{N}{2}$. The fact that $n_{i}|N_{i-1}$ and $n_{i}|\frac{N}{2}$, ensures that $n_{i}|\frac{N}{2}-N_{i-1}$. This implies in turn that $\frac{N}{2}-N_{i-1}\geq n_{i}$ and thus $N_{i}\leq\frac{N}{2}$square
@@ 4 The CHS chaotification model used for diffusion
We will run the dynamical system with a seed depending on the total intensity scaled between 00 and $1$ of the set of images to be encrypted. A total intensity of 00 corresponds to all white images, while a total intensity of $1$ corresponds to all black images. Thus, we set
$x_{0}:=\sum_{\begin{array}[]{l}\qquad\qquad 0\leq x\leq 2^{n}-1\\ \qquad\qquad 0\leq y\leq 2^{n}-1\\ \qquad\qquad 0\leq m\leq M^{{}^{\prime}}-1\end{array}}\frac{C_{xym}}{M^{{}^{\prime}}(2^{L}-1)2^{2n}},$
@@
$\lfloor T_{s_{m,i}}(y_{m,2^{n}-i+1})T_{t_{m,j}}(x_{m,2^{n}-j+1})\times 10^{q_{m}}\rfloor\;\text{mod}\,2^{\lceil log_{2}M\rceil},$
with for instance $q_{m}\geq 4$. From there, obtain a $\lceil log_{2}M\rceil$-qubit secret key
$|K_{\lceil log_{2}M\rceil-1,m,i,j}\dots K_{0,m,i,j}>.$
@@ 6 An example of quantum implementation for the scrambling
[…]
The discrete baker map $B_{(4,2,2)}$ is defined on the integer lattice $[0,2^{3}-1]\times[0,2^{3}-1]$ by:
@@
We will now tackle the quantum circuit for $B=B_{(2^{4},2^{3},2^{3},2^{5},2^{6},2^{7})}$. By using Lemma $1$ of § $\,3$, we are able to compute the subfunctions like follows:
@@
$B(x,y)=\begin{cases}f(x,y)&\text{when $x_{4}=x_{5}=x_{6}=x_{7}=0$}\\ (g\circ f)(x,y)&\text{when $x_{7}=1$}\\ (h\circ f)(x,y)&\text{when $x_{6}=1\,\&\,x_{7}=0$}\\ (k\circ f)(x,y)&\text{when $x_{5}=1\,\&\,x_{6}=x_{7}=0$}\\ (l\circ f)(x,y)&\text{when $x_{4}=1\,\&\,x_{5}=x_{6}=x_{7}=0$}\\ \end{cases}$

### 22 Mar 2026 v2

@@
Abstract
Due to the strong correlations among adjacent pixels, bulky data volume, and the handling of various data formatting, traditional encryption schemes are not suitable for image encryption. After introducing a bit-plane quantum representation for a multi-image, we present a novel way to encrypt/decrypt multiple images using a quantum computer. Our encryption scheme is based on a two-stage scrambling of the images and of the bit planes on one hand and of the pixel positions on the other hand, each time using quantum baker maps. The resulting quantum multi-image is then diffused with controlled CNOT gates using a sine chaotification of a two-dimensional Hénon map as well as Chebyshev polynomials. The decryption is processed by operating all the inverse quantum gates in the reverse order. All the secret keys that are used for the quantum encryption process have been previously sent from the sender to the receiver by post-quantum cryptography. The cryptographic scheme can also be used on a classical computer.
@@ 1 Introduction
Chaotic systems are commonly used as diffusion tools in quantum image encryption methods. The idea to use chaos in encryption algorithms originates in [13]. Chaotic functions have the characteristics of generating completely different sequences through relatively small changes in the tuning parameters or in the initial conditions. One of the best-studied such functions is the logistic function which has been used to model biological populations. However, the parameter range for chaotic behavior is relatively small and for some isolated parameter values, the behavior is not chaotic. A general way to enlarge the parameter range of chaos and in fact make this range infinite is to use a sine chaotification model composed with another given chaotic map, see [7]. This idea was used by the authors in [11], applied to the $1D$ logistic map. In the present paper we also apply this idea with the base chaotic map being the $2D$ Hénon map. For each scrambled image, we use a different control parameter, while the two initial conditions depend on the multi-image plaintext. For a given scrambled image, the two pseudo-random sequences obtained, of length the side size of each image, are sorted by increasing order (we impose the elements of each sequence to be distinct with one another). The resulting two pseudo-permutations are then used in the following manner. Each element of the first (resp second) pseudo-random sequence gets applied a Chebyshev polynomial with index issued from the second (resp first) pseudo-random permutation and the resulting two values get multiplied together yielding an element of the interval [1,1][-1,1]. For each pixel position we thus obtain an element of this interval. We then multiply it by a power of ten depending on the image and take the integer part modulo the $\lceil log_{2}L\rceil$-th power of two, where $L$ is the number of bit planes. Each integer is then converted into a sequence of bits of length $L$, namely the secret keys, that are ready to be XOR-ed with the bit values of each bit plane.
@@ 2 The BRQMI: a bit-plane representation for quantum multiple images
where $P_{lxy}$ is the bit value in the $l$-th bit plane at pixel $(x,y)$. We now introduce the new model BRQMI used for our encryption/decryption scheme. Our BRQMI is a combination of BRQI and of the quantum representation model for multiple images QRMMI of [22]. With respect to BRQI, $\lceil log_{2}M\rceil$ qubits are added in order to encode the $m$-th image, with $0\leq m\leq M-1$ and $M$ the total number of images to be encrypted. Denoting by $L$ the number of bit planes, the quantum image is the superposition:
@@
$P_{mlxy}$ is the bit value in the $l$-th bit plane of the $m$-th image at pixel $(x,y)$. The 0-th bit plane denotes the least significant bit plane and the (L1)(L-1)-th bit plane is the most significant bit plane. With this representation, it is understood that we set:
@@
Moreover, for the sake of our encryption/decryption scheme, because as part of the two stage scrambling process we will apply a quantum baker map in the $(m,l)$ plane, we impose that $M=L$. This will be assumed throughout the rest of the paper. As far as the conversion of the classical multiple images into the quantum multiple images, it does translate in the following way. On one hand, if $M>L$, that is there are more images than bit planes, then we add more bit planes filled with bit values 0. On the other hand, if $M<L$, that is there are less images than bit planes, then we add LML-M blank images. It has to be determined in both cases how this would affect the security of the protocol. We note in particular that in the case when there are more bit planes than images, the attacker has no way of deciding on the number of images that are encrypted. This is rather a factor increasing the security of the scheme.
@@ 3 The quantum baker map used for scrambling
$M_{s}(x,y)=\Bigg(2^{n-s}x\,\text{mod}\,2^{n}+y\,\text{mod}\,2^{n-s},\frac{y-y\,\text{mod}\,2^{n-s}}{2^{n-s}}+x-x\,\text{mod}\,2^{s}\Bigg)$
@@
Proof. Suppose $N_{i-1}\leq\frac{N}{2}$. The fact that $n_{i}|N_{i-1}$ and $n_{i}|\frac{N}{2}$, ensures that $n_{i}|\frac{N}{2}-N_{i-1}$. This implies in turn that $\frac{N}{2}-N_{i-1}\geq n_{i}$ and thus $N_{i}\leq\frac{N}{2}$ $\square$
@@ 4 The CHS chaotification model used for diffusion
We will run the dynamical system with a seed depending on the total intensity scaled between 0 and $1$ of the set of images to be encrypted. A total intensity of 0 corresponds to all white images, while a total intensity of $1$ corresponds to all black images. Thus, we set
$x_{0}:=\sum_{\begin{array}[]{l}\hskip 16.38895pt\hskip 16.38895pt0\leq x\leq 2^{n}-1\\ \hskip 16.38895pt\hskip 16.38895pt0\leq y\leq 2^{n}-1\\ \hskip 16.38895pt\hskip 16.38895pt0\leq m\leq M^{{}^{\prime}}-1\end{array}}\frac{C_{xym}}{M^{{}^{\prime}}(2^{L}-1)2^{2n}},$
@@
$\lfloor T_{s_{m,i}}(y_{m,2^{n}-i+1})T_{t_{m,j}}(x_{m,2^{n}-j+1})\times 10^{q_{m}}\rfloor\;\text{mod}\,2^{2^{\lceil log_{2}M\rceil}},$
with for instance $q_{m}>M$. From there, obtain a $2^{\lceil log_{2}M\rceil}$-qubit secret key
$|K_{2^{\lceil log_{2}M\rceil}-1,m,i,j}\dots K_{0,m,i,j}>.$
@@ 6 An example of quantum implementation for the scrambling
Suppose the size of the image is $256\times 256$. By the theorem of § $\S\,3$, there are $P_{8}=1947270476915296449559703445493848930452791205$ admissible partitions for the quantum baker map, one of which is $(16,8,8,32,64,128)$. This is the partition we chose for our example of quantum implementations of pixel positions scrambling. Suppose also that there are $8$ images to encrypt at once and the gray values are spread over $8$ bit planes. Amongst the $P_{3}=26$ admissible partitions for the bit planes and images mixing, we chose to study the quantum circuit for $(4,2,2)$.
The discrete baker map $B_{(4,2,2)}$ is defined on the integer lattice $0,2^{3}-1]\times[0,2^{3}-1]$ by:
@@
We will now tackle the quantum circuit for $B=B_{(2^{4},2^{3},2^{3},2^{5},2^{6},2^{7})}$. By using Lemma $1$ of § $\S\,3$, we are able to compute the subfunctions like follows:
@@
[…]